# Designs for Enabling Collaboration in Human-Machine Teaming via Interactive and Explainable Systems

### v1

@@ 5 Human-Subjects Study
We describe the reward scheme and domains further in the appendix. Importantly, we assess scenarios in which the team is rewarded more for collaboration than teammate-agnostic or independent behavior, which represents the highest level of teamwork Kolbeinsson et al. (2019).
(a) Forced Coordination
(b) Optional Collaboration
Figure 4: This figure shows the two domains used in our experiment.
For each domain, we train an IDCT policy via agent-agent collaborative training (for conditions IV1:C1-4) and a NN policy following the population-based training scheme in Strouse et al. (for condition IV1:C5). In IV2-D1, the IDCT policy (utilized in IV1:C1-4) converged to a policy with an average reward of $315.22\pm 14.59$, and the neural network policy converged to an average reward of $403.16\pm 16.08$ over 50 teaming simulations with the synthetic human teammate the policy was trained with. In IV2-D2, the IDCT policy converged to a policy with an average reward of $171.46\pm 18.89$, and the neural network policy converged to an average reward of $295.02\pm 1.86$. Thus, a consequent confound due to the current difference in performance capabilities between interpretable vs. black-box models is that the NN policy outperforms the IDCT policy in both domains. This displays a need for improving optimization algorithms for interpretable models. However importantly, while the initial simulated performance of interpretable models may underperform black-box models, the ability for humans to understand machine behavior and improve upon behavior may allow these approaches to compete or even outperform black-box NN models. We can also compare to the heuristic policies presented in Section 3, observing that the training performance of the IDCT and NN policies in IV2-D2 underperform the collaborative heuristic (408 vs. 295.02 and 171.46). We provide the trained IDCT policies for each domain in the appendix, finding that after pruning, the AI policy has two and three leaves, respectively.
Procedure: A participant is first randomly placed into one of the five conditions in IV1. The participant starts with a pre-experiment survey collecting demographic information, experience with video games and decision trees, and the Big Five Personality Questionnaire Chmielewski and Morgan (2013). Afterward, a participant conducts a brief tutorial in Overcooked with a random AI agent, improving the user’s understanding of game controls and the assigned task. Once completed, the primary experimentation begins. Users will team with an AI four times in each domain (randomly ordered), and are told that their goal is to maximize their score in the last teaming interaction, the “performance round.” After each teaming interaction, in the first three factors, the user will modify and visualize the AI’s policy (IV1-C1), the AI will optimize its own policy proceeded by user visualization (IV1-C2), or the user will solely view the policy (IV1-C3). In IV1-C4 and IV1-C5, as the AI is black-box (perceived to be black-box in IV1-C4 and truly black-box in IV1-C5), transitionary pages are shown to the participant, providing them a pause before they team with the agent again. Upon completion of the condition-specific (or lack of) actions, users complete a NASA-TLX Workload Survey. After users have completed a domain, providing us with four episodes of teaming data and workload assessments, we administer several post-study scales, including the Human-Robot Collaborative Fluency Assessment Hoffman (2019), Inclusion of Other in the Self scale Aron et al. (1991), and Godspeed Questionnaire Bartneck et al. (2019). Upon completion of the two domains, the experiment concludes. By comparing conditions, we gain insights into the advantages of interpretability and interaction.
(a) Performance Data in IV2-D1: Forced Coordination
(b) Performance Data in IV2-D2: Optional Collaboration
Figure 5: Gameplay scores from users across teaming iterations.
5.1 Results
Our experiment is a 5 (teaming method; between-subjects) $\times$ 2 (no. of domains; within-subjects) $\times$ 4 (no. of repeated evaluations; within-subjects) mixed-factorial experiment. We recruited 50 participants under an IRB-approved protocol, whose ages range from 18 to 32 (Mean age: 24.14; Std. Dev.: 4.10; 46% Female, 52% Male, 2% Non-Binary), with participants randomly assigned to each of the factor levels, with ten total subjects per level. The duration of the experiment was $70.98\pm 19.71$ minutes. Our data was modeled as a full-factorial, between-subjects ANOVA. We test for normality and homoschedascity and employed a corresponding non-parametric test if the data failed to meet these assumptions. We display our objective findings in Figure 5.
RQ1: Team Coordination Performance: In analyzing team reward, we find trends with respect to the maximum reward participants obtained within a domain across all teaming iterations (Figure 6). Utilizing a Friedman’s test, we find that there is a significant difference across domains ($\chi^{2}$(1)=46.08, $p<0.001$). Accordingly, we analyze the domains separately.
Figure 6: Maximum Reward Across Each Condition (top) and Subjective Ratings for Positive Traits and Team Fluency (bottom).
In IV2-D2, a Kruskal-Wallis Test was conducted to analyze differences in participant teaming performance across conditions, finding a significant effect ($\chi^{2}(4)=29.922,p<0.001$). We conduct post-hoc pairwise comparisons, utilizing Dunn’s test, and find that IV1-C5 (Ficticious Co-Play) is significantly better than IV1-C2 ($p<0.001$), IV1-C3 ($p<0.001$), and IV1-C4 ($p<0.001$), and IV1-C1 (Human-Led Policy Modication) is significantly better than IV1-C2 ($p<0.05$), IV1-C3 ($p<0.05$), and IV1-C4 ($p<0.05$). For white-box AI teammates (IV1:C1-3), the latter finding displays the benefit of Human-Led Policy Modification in improving HMT performance for interpretable models.
These findings display that 1) white-box approaches supported with policy modification can outperform white-box approaches alone, 2) black-box models can outperform white-box approaches in HMT, and 3) by comparing IV1-C3 to IV1-C4, interpretability afforded via tree visualizations did not provide any direct objective benefits. Finally, in Optional Collaboration, across all conditions we see that HMT scores are not near that of the collaborative heuristic, displaying a gap that must be addressed to achieve effective HMT.
RQ2: Team Development: In analyzing RQ2, we look at the change in reward across iterations one to four and relate our findings to Tuckman’s model. Utilizing a Friedman’s test, we find that there is a difference across domains ($\chi^{2}$(1)=20.48, p$<$0.001). Thus, we analyze the domains separately.
We conduct an analysis to see which conditions facilitate team development. In IV2-D1, we see that none of the conditions results in a significant improvement in teaming performance over repeated iterations. In IV2-D2, we see IV1-C1 ($p<0.01$) and IV1-C2 ($p<0.01$) significantly improve over repeated teaming interactions. The improving interactions can be connected to the Norming stage in team development, where teams begin to develop a strategy and shared mental models. We see that conditions that facilitate Norming have the attribute of policy adaptation and are white-box.
Next, we analyze whether different person-specific factors allow HMT to improve more quickly than others. In IV2-D1, we find that conscientiousness is trending in its correlation with faster improvement ($0.10.05<p<$ 0.1). In IV2-D2, we find that there is a significant effect in how much participants improve and their familiarity with Trees ($F(1)=7.448,p<0.01$). These findings signify that positive interaction with interpretable models may be more beneficial to those with an engineering background and specific personality traits.
Finally, we detect an interesting trend in IV2-D1 under the IV1-C1 condition. We see a drop in performance between the first teaming iteration and later iterations, followed by a rising trend. We believe this relates to the Forming and Storming stages, where team members are still developing effective strategies to coordinate. As the last iteration shows an improvement in performance, we hypothesize that the team was shifting into the Norming stage. In future, it would be interesting to evaluate a larger number of teaming iterations to see if the behavior would continue to trend upward. This is an additional challenge that requires further research as more iterations require more experimentation time and resources.
Subjective Findings: In IV2-D1, we find that users did not find any subjective differences toward the teaming interaction across conditions. In IV2-D2 (Figure 6), we find that users find collaboration with AIs under condition IV1-C2 and IV1-C4, on average as less fluent than IV1-C1 (p$<$0.01, p$<$0.01), and IV1-C4 as less fluent than IV1-C5 ($p<0.05$). Users also trusted the AI and perceived the AI contributed more in IV1-C5 than in IV1-C2 (p$<$0.05, p$<$0.05) and IV1-C4 ($p<0.05,p<0.05$). Furthermore, the users viewed the AI more positively in IV1-C1 and IV1-C5 than in both IV1-C2 (p$<$0.05, p$<$0.05) and IV1-C4 (p$<$0.05, p$<$0.01). As seen through these findings, the ability to interact with an interpretable model is perceived significantly better across several measures. We provide our codebase for our experiment here.
Design Guidelines: Achieving fluent HMT is challenging. Toward this goal, we specify the following guidelines.

### v2

@@ 5.1 Results
    The design of learning schemes to support the generation of collaborative AI behaviors rather than individual coordination. We need techniques that avoid converging to the local maxima of individual coordination and scenarios that allow for properly evaluating cooperation.
  3. 3.
    The creation of mixed-initiative interfaces that enable users, who may vary in ability and experience, to improve team collaboration across and within interactions. As we found a large diversity in perceived usability of our interface (finding an average score of $58.25\pm 27.61$, with some users finding the interface good ($>$75) and others poor ($<$35)), effective interfaces are vital in shifting from only a subset of users benefiting to all users being able to create effective teammates.
  4. 4.
    The evaluation of teaming in a larger number of interactions. As agents are deployed, team performance will change over time, going through a transient period before reaching peak performance. Understanding this process of team development is essential in creating high-performance HMT.
6 Conclusion
This work investigates repeated interactions with machine learning models within a sequential decision-making HMT paradigm. We present a key gap in HMT, displaying that current methods do not facilitate human-machine collaboration to the fullest. We find that human-led policy modification allows for a team to achieve higher performance than white-box models without this capability. However, as interpretable models are more difficult to generate, Fictitious Co-Play is able to better support high performance. Given these mixed findings, future work must focus on developing better white-box teammates, study the modality of communication in HMT, and explore mechanisms to allow HMT to scale beyond individual coordination and toward effective collaboration.
Appendix A Appendix
In the Appendix, we provide further information regarding our testbed for Human-Machine Collaboration, Overcooked-AI (Section B), additional model and training details for our interpretableML architecture, the Interpretable Discrete Control Tree (Section C), complete information regarding our statistical analysis (Section E), further discussion regarding our paper’s results, limitations, and future work (Section F), and finally, a working definition of what we mean by “interpretable".
Appendix B Overcooked-AI
Overcooked-AI [5] is a testbed to evaluate human-AI interaction and has been used across numerous prior work studying human-AI collaboration [40, 10]. Here, two agents are tasked with creating and delivering as many soups as possible within a given time. Achieving a high score requires agents to navigate a kitchen and repeatedly complete a set of sequential high-level actions, including collecting ingredients, placing ingredients in pots, cooking ingredients into a soup, collecting a dish, getting the soup, and delivering it. Both players receive the same score increase upon delivering the soup. We modify the original Overcooked-AI game to be a simultaneous-move game as opposed to the original formulation of allowing agents to perform actions asynchronously. This modification prevents the collaborative score metric from being dominated by super-human AI speed, causing the overall score to be more reliant upon effective collaboration and strategy.
We utilize two map configurations we term, Forced Coordination and Optional Collaboration, displayed in Figure 4 of the main paper. Each domain was chosen so that collaborating with the teammate would result in a higher score than working individually. In our newly-created domain, Optional Collaboration, creating mixed-ingredient dishes (combining onions and tomatoes) will receive a higher score than single-ingredient dishes. Teammates have 200 timesteps to collaborate and cook as many dishes as possible.
State-Space, Action-Space, and Reward Scheme
Policies reason over a semantically meaningful 13-dimensional feature space as opposed to pixel space, detailing the objects each agent is holding, pot statuses, and counter objects. Each of these features is binary. For the action space, instead of using cardinal actions, we allow the AI to utilize macro-actions that can accomplish high-level objectives such as ingredient collection, ingredient placement, and soup serving. Macro-actions are planned using an A* planner, and we perform dynamic replanning at each timestep. Prior work has shown macro-actions can enhance interpretability [4]. This state and action space allow forlearning an interpretable tree-based policy that can be understood and manipulated by end-users.
In Forced Coordination, for the reward scheme, we follow a similar distribution as prior work and give a reward score of 60 per dish served, 3 for an item placed into a pot, 3 for a useful dish pickup, and 5 for a soup pickup. In Optional Collaboration, for the reward scheme, we give a reward score of 50 for a mixed-ingredient dish, 30 for a single ingredient dish, 3 for an item placed into a pot, 3 for a useful dish pickup, and 5 for a soup pickup.
Appendix C Additional IDCT Model Details
Here, we provide additional model details for the proposed Interpretable Discrete Control Tree (IDCT).
C.1 Architecture
Our IDCTs are based on differentiable decision trees (DDTs) [41] – a neural network architecture that takes the topology of a decision tree (DT). DDTs contain decision nodes and leaf nodes; however, each decision node within the DDT utilizes a sigmoid activation function (i.e., a “soft" decision) instead of a Boolean decision (i.e., a “hard" decision). Each decision node, $i$, is represented by a sigmoid function, displayed as $y_{i}=\frac{1}{1+\exp(-\alpha(\vec{w}_{i}^{T}\vec{x}-b_{i}))}$, where $\vec{w}_{i}$ and $b_{i}$ represents the weight and bias terms of the decision node, respectively. As this representation is difficult to interpret, [32] presented differentiable crispification, consisting of two components: 1) Decision node crispification, which recasts each decision node to split upon a single dimension of our input feature, and 2) Decision outcome crispification, which translates the outcome of a decision node so that the outcome is a Boolean decision rather than a set of probabilities. Both operations utilize the straight-through trick [3] to maintain gradients, allowing for both an interpretable forward propagation through the model that traces down a single branch of a tree as well as gradient flow to update parameters of the neural tree model. We utilize this approach in our IDCTs to maintain interpretability.
We initialize our IDCTs to be symmetric complete decision trees with $N_{l}$ decision leaves and $N_{l}-1$ decision nodes. Each decision leaf is represented by a sparse categorical probability distribution over actions. At each timestep, a state variable is propagated through each decision node, split on a single decision rule, with the output being a Boolean causing the decision to proceed via the left or right branch until arrival at a leaf node. At each leaf node, we sample from the respective probability distribution to produce a macro-action (e.g., “get an onion" or “place held ingredient on counter").
C.2 Training
For training this model, we utilize agent-agent collaborative training where an interpretable tree-based agent (maintaining an IDCT) is paired with a second policy (representing the human player), and both models are trained via decentralized PPO [39]. It is important to note that each agent maintains its own buffer and optimizers. Further, we improve model predictability by applying an L1 norm loss over leaf node distributions for the IDCT agent to ensure sparsity, penalizing high entropy action distributions at a leaf. Our training procedure mimics that of PPO, utilizing a modified loss function displayed in Equation 1, and policy update in Equation 2, where $\theta$ represents the aggregate set of weights for the IDCT, $\hat{A}_{t}$ represents the advantage estimate at time $t$, and $a_{l}$ represents the distribution maintained at each leaf, $l$.
$\begin{split}L(\theta)&=\mathbb{E}_{\tau}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\right]\\ &+\sum_{1}^{L}\lambda|a_{l}|\end{split}$ (1)
$\theta_{k+1}=\operatorname*{arg\,max}_{\theta}L(\theta)$ (2)
C.3 Contextual Pruning
As we focus on creating agents that cooperate with humans, we must limit the size of our interpretable tree-based models to a certain depth to promote user understanding. This follows prior work, finding trees of arbitrarily large depths can be difficult to understand [13] and simulate [24], and that a sufficiently sparse DT is desirable and considered interpretable [20]. However, this can make training difficult, as a small tree may not have the representational power to learn a high-performing policy.
Algorithm 1 Contextual Pruning Algorithm
Input: IDCT I(.)
Output: Pruned IDCT
1:  Set__\__Node__\__Domains(IDCT=I, minValue=0, maxValue=1)
2:  queue = [I.root]
3:  while queue is not empty do
4:     currentNode $\leftarrow$ queue.pop()
5:     if currentNode.compareValue $<$ currentNode.lowerBound then
6:        currentNode.prunable = True
7:     end if
8:     if currentNode.compareValue $>$ currentNode.upperBound then
9:        currentNode.prunable = True
10:     end if
11:     Update__\__Domains__\__For__\__Children(currentNode, lowerBound, upperBound, currentNode.compareValue)
12:     Add__\__Children__\__To__\__Queue(currentNode, queue)
13:  end while
14:  I $\leftarrow$ Prune__\__Nodes__\__From__\__Tree(I)
15:  return  I
In Algorithm 1, we present details of how contextual pruning is accomplished. In Step 1, we initialize a domain vector representing the current minimum and maximum values for each feature. Since our Overcooked domain utilizes binary features, all bounds are initialized to 0 and 1. Formally, this can be written as by the Cartesian product $B=[0,1]\times\cdots[0,1]$, of cardinality $d$ (where $d$ is the dimensionality of the state space). In Step 2, we initialize a queue that will be used to perform a breadth-first search to visit each node in a hierarchical order. In Step 4, we receive a node from the queue. In Step 5, we check the threshold value of the current node and compare it to the current node’s vector of minimum values. This operation looks to see if the node results in a tree sub-space that is out of bounds (i.e., impossible to reach). We perform a similar computation in step 8, checking the maximum values. In both cases, we look to find child nodes that do not yield a reduction in the hyperspace as candidates for pruning. In Step 11, we update the children based on the threshold value of our current node and its sign (as we can have $<$ or $>$ within a node), creating a new bounding box. In step 12, we add the children of the current node to the queue, and loop back to Step 4, repeating steps 5-12 until the queue is empty. In Step 14, we prune tree sub-spaces that are impossible to reach.
C.3.1 Computational Analysis
The computational complexity of our contextual pruning algorithm can be analyzed in terms of both time and space complexity. In terms of time complexity, it is equivalent to that of Breadth-First Search (BFS), specifically, $\mathcal{O}(V+E)$, where V denotes the number of vertices and E represents the number of edges in the tree. Regarding space complexity, our algorithm exhibits similar characteristics to BFS for trees with only two leaves. In such cases, the space complexity of BFS is $\mathcal{O}(V)$, as it stores all the vertices at the maximum breadth level in the queue during the traversal. Consequently, the space complexity of our contextual pruning algorithm is also $\mathcal{O}(V)$, making it efficient and scalable for trees with a limited number of leaves.
Utilizing contextual pruning alongside our training framework allows us the benefit of training large tree-based models, greatly improving ease-of-training, while still being able to simplify the resultant model to a smaller, equivalent representation.
C.3.2 Results of Pruning
To evaluate the utility of pruning, we train models of various sizes (8-leaf, 16-leaf, 32-leaf, 64-leaf, 128-leaf, 256-leaf) in Forced Coordination and perform pruning on the resultant model. We find that models of larger size converge to higher performance (i.e., easier-to-train), following prior work displaying the utility of larger models. Further, empirically, we find we can reduce model sizes by 64-128x in tree depth. We provide a pipeline to allow for model training and contextual pruning in our GitHub repository https://github.com/CORE-Robotics-Lab/Team-Development-with-Transparent-Policies.
C.4 Hyperparameters
In IV2-D1: Forced Coordination and IV2-D2, we train an IDCT with 256 leaves, a learning rate of $1e^{-3}$, and regularization parameter of $1e^{-4}$. This hyperparameters were chosen through trial and error, where we find larger models with a small learning rate and regularization exhibited greater learning early on. The rest of the parameters follow default parameters from the PantheonRL codebase [38] for training Overcooked agents. After contextual pruning, in both domains, we end up with an AI policy with two and three leaves in Forced Coordination and Optional Collaboration, respectively.
For training fictitious co-play agents, we train 32 models of teammates in each domain, saving policies at every 100 epochs. At the end of training, we sort the performance of saved policies and utilize the initial, mid-performing, and highest to create our population of diverse agents, totaling 96 agents. A neural network model is then paired in a multi-task training framework to team with this agent.
Our models are all trained on a local desktop computer containing a Nvidia RTX 2080 GPU and 16 GB of CPU memory. Training time for each agent took approximately 12 hours across a single core. We provide further instructions to replicate our models within the above codebase.
We include a high-level diagram of how IDCT agents are generated in Figure 6.
Figure 6: Tree Policy Generation for Conditions IV1-C1-C4
C.5 Visualization of IDCT Policies in Each Domain
Here, we present visualizations of trained IDCT models in each domain. As seen in Figures 7 and 8, the resultant policies have two and three leaves for the Forced Coordination and Optional Collaboration domains, respectively. Note that these images are pulled from our interface and thus have extra annotations to improve readability.
Figure 7: Trained Interpretable Discrete Control Tree in the Forced Coordination Domain.
Figure 8: Trained Interpretable Discrete Control Tree in the Optional Collaboration Domain.
Appendix D Additional User Study Information
Our experiment was reviewed and approved by the Institutional Review Board at the Georgia Institute of Technology under Protocol Number H23043. All participants in our experiment signed a consent form, received a description of the risks involved in our study, and received compensation for participating. Below, we describe specifics regarding the consent procedure, additional details that describe the experiment procedure, and the compensation scheme.
D.1 Consent Procedure
At the start of the experiment, the participant is provided a consent document. This document describes the purpose of the experiment, exclusion/inclusion criteria, the experiment procedure, the risks of the experiment, the compensation scheme, and details regarding data storage and confidentiality.
D.2 Additional Information Regarding Specific Conditions
IV1-C1: Human-Led Policy Modification is enabled through the contribution of the interpretable machine learning architecture to train collaborative AI teammates, a training advancement to enhance interpretability, and a mechanism to allow humans to modify the tree in simple ways, including tree deepening, decision variable modification, and leaf node modification. The following conditions: IV1-C2: AI-Led Policy Modification, IV1-C3: Static Policy - Interpretability, and IV1-C4: Static Policy - Black-Box all utilize the same architecture and starting policy but ablate different components of the interaction and interpretability.
After a teaming episode in the IV1-C2: AI-Led Policy Modification condition, the AI utilizes recent gameplay to fine-tune a human gameplay model via Behavioral Cloning and performs reinforcement learning for five minutes to optimize its own policy to better support the human teammate. In this collaborative agent policy optimization stage, we utilize the parameters described in Section C.4 and add a timer to stop the optimization. Upon completion of policy optimization, we check if the policy has improved through simulated interactions with the behavior cloning agent, and if so, update the policy. In the case that the policy degrades, we use the original policy prior to optimization. The user can visualize the updated AI policy in its interpretable tree form prior to the next teaming interaction.
IV1-C3: Static Policy - Interpretability and IV1-C4: Static Policy - Black-Box are static policies that do not change across repeated gameplay. Thus, we do not have any specific additional hyperparameters to discuss within the appendix.
To improve the transparency of the conditions in our experiment, we provide a flow diagram that displays the interaction being assumed within each condition in Figure 9.
(a) Experiment Flow for IV1-C1: Human-Led Policy Modification
(b) Experiment Flow for IV1-C2: AI-Led Policy Modification
(c) Experiment Flow for IV1-C3: Static Policy - Interpretability
(d) Experiment Flow for IV1-C4: Static Policy - Black-Box and IV1-C5: Fictitious Co-Play
Figure 9: This figure displays an experiment flow diagram for each condition.
D.3 Compensation Scheme
Participants were compensated at a rate of 20 US dollars per hour of the experiment.
Appendix E Complete Statistical Analysis
Here, we present complete details regarding our analysis, including all test statistics as well as nonsignificant and trending comparisons.
E.1 RQ1: Team Coordination Performance
As mentioned in the main paper, we allow humans to team with the AI across four episodes, providing us with four teaming scores. Within the main paper, we reported differences with respect to the maximum score participants were able to obtain across iterations. Here, we analyze data in the performance round (the last iteration), where participants were told to maximize performance. We note that participants self-reported their gaming familiarity (100-point scale) and weekly hours playing video games. Across all participants, self-reported gaming familiarity was rated as $73.19\pm 23.80$ and weekly gaming hours was $4.44\pm 5.32$. This information was used in our statistical analysis, and significance was not found in performance variation as a function of gaming expertise. Utilizing a Friedman’s test, we find that there is a significant difference across domains ($\chi^{2}(1)=38.7,p<0.001$). Accordingly, we analyze the two domains separately.
In IV2-D1, we find our data does not meet the necessary assumptions and utilize non-parametric tests. A Kruskal-Wallis Test was conducted to analyze differences in performance round reward across conditions, and we find a significant effect ($\chi^{2}(4)=20.85,p<0.001$) across conditions. We conduct post-hoc pairwise comparisons, utilizing Dunn’s test, and find that IV1-C5 is significantly better than IV1-C1 ($p<0.01)$, IV1-C3 ($p<0.01)$, and IV1-C4 ($p<0.01$). IV1-C5 is trending as significantly better than IV1-C2 with a p-value of 0.0275 (significance is $<0.025$ or $(\alpha/2)$ due to the Bejamini-Hochberg adjustment).
[…]
These results are similar to those in the paper when analyzing the maximum reward and result in a similar set of conclusions: 1) black-box models can outperform white-box approaches, and 2) white-box approaches with policy modification have some benefit over white-box approaches alone. Further, as we see that tree familiarity positively correlates with performance round rewards, exploring alternative paradigms, such as natural language for describing and programming trees may benefit users unfamiliar with decision trees.
E.2 Team Development
Here, we analyze the trends across iterations (did agents improve from iteration one to four) and identify characteristics of users that performed well in team development. Utilizing a Friedman’s test, we find that there is a significant difference across domains ($\chi^{2}$(1)=20.48, $p<0.001$).
We conduct separate Wilcoxin signed-rank tests for each condition, and utilize the Bonferroni correction in determining significance ($\alpha/5$). In IV2-D1, we see no condition significantly improves significantly over repeated iterations. In IV2-D2, we find that IV1-C1 ($p<0.01$) and IV1-C2 ($p<0.01$) significantly improve over repeated teaming interactions.
Appendix F Discussion, Limitations, Future Work, and Societal Impacts
Discussion: In this paper, we provide several contributions towards interactive HMT. We first present weaknesses in prior work, displaying that learned collaborative agents can be individualistic and rigid. To address these weaknesses, we propose an interactive scheme termed human-led policy modification to bridge the gap between individualized coordination and adaptive, effective collaboration. We do so by creating a feedback loop that facilitates team policy changes during HMT. This is accomplished by 1) utilizing an interpretable policy representation to provide users with insight into the teammate’s decision-making, specifically the IDCT, an interpretable tree-based model that can be trained via reinforcement learning and pruned to a smaller, equivalent representation, and 2) creating a user interface to support the end-user modifying the policy to their evolving specifications. We deploy and compare our interactive policy modification scheme to several other techniques, including two popular prior works and variations of our proposed condition. While we do not a direct objective benefit of human-led policy modification compared to utilizing a black-box model supported with a population-based training scheme [40], we find important takeaways that motivate the importance of conducting longer-term, repeated-interaction studies. Specifically, white-box approaches that facilitate interpretation can be used within a feedback loop to lead to policy improvement, users may require a larger number of interactions to reach a team consensus and maximal performance, and there are person-specific characteristics that may lead to some users being able to take advantage of interpretable models and interaction more than others.
Limitations: This study was conducted at a university. While the population was diverse in age, gender, and university major, all students had some college education and most students were based in engineering, presenting a population bias. Furthermore, the population represented by the age group of 18 to 32 years old (mean of 24.14, std of 4.1) within our experiment may not directly generalize to an older population with extensive training. Furthermore, the experiment findings may not generalize to all contexts and scenarios within HMT. We reiterate that our findings are within a two-agent human-machine team within a relatively low-dimensional and short-horizon game, Overcooked-AI. In scaling to more complex and dynamic environments, the tree size needed to represent a high-performing agent will likely increase. In these cases, users may require more time to interact with and understand an agent’s policy. There may be several capabilities that can be added to the Human-Led Policy Modification interaction paradigm, which may make the process quicker and easier. For example, model verification or forward simulation can be used to provide the human with other types of feedback prior to the next teaming iteration. Furthermore, for increasingly complex games, agent policies can also operate over different levels of abstraction, providing the human with a tradeoff with fine-grained control of the agent policy and tree size. Finally, different policy visualizations may better support certain populations of users, emphasizing the need for collecting user background information and future research in interpretability for embodied agents.
Future Work: In the future, it would be interesting to conduct a similar experiment to a higher number of iterations, or until the team converges to a set of coordination strategies (the “performing" stage in Tuckman’s model). Further, the possibility of adding in feedback from the AI regarding human-led policy modification (checking for logic inconsistencies, etc.) may be used to facilitate speedier team development. It would also be interesting to utilize different paradigms in communicating with the human as language may be an easier medium than a decision tree interface. Future work should also be done to optimize the accessibility of GUIs for policy modification via xAI techniques. Finally, expanding this research to real-world collaborative robot settings in healthcare of manufacturing that utilize tree-based policies, such as collaborative packaging [12, 17] or agile robotics [8, 21], would lead to additional insight into human-machine team development with robot teammates.
Positive and Negative Societal Impact: This work investigates repeated interactions with interpretable machine-learning-based agents in a collaborative game. As autonomous agents (e.g., robots) are deployed in the real world, insights from this work may be applied to assist in creating a fruitful working relationship between a human and an agent. We do not believe this work has any negative societal impacts.
Appendix G Working Definition of Interpretability
As mentioned in the main paper, our agent representation is that of an Interpretable Discrete Control Tree, which reasons over a state space with high-level binary features and multi-step macro-actions. This model (which, in layman’s terms, is a decision tree with action probabilities at each node) is the true learned model produced via reinforcement learning, not an abstraction created post hoc. This model is interpretable as its representation is “constrained in model form so that it is either useful to someone, or obeys structural knowledge of the domain, such as monotonicity, causality, structural (generative) constraints, additivity, or physical constraints that come from domain knowledge" [35]. In our case, the model constraints are inherent within the novel IDCT architecture, and the utility of this model to a user is that this model 1) is able to provide users with some awareness over the agent’s behavior (and possibly, simulate the agent’s decision making) and 2) provides users with the ability to explicitly modify agent behavior (a capability not possible with black-box models).
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.